1. If a car is moving at a speed of V, the stopping distance is D. If the car is moving at 2V, what
would be the stopping distance? (Please explain)
- D
- 2D
- 3D
- 4D
- 5D

2. The temperature of a transformer increases is/are due to
- Ohmic loss,
- Hysteresis loss,
- eddy current

3. An object measures 1m when it is restless. What speed you need to move in order to
measure its length to be 0.8m?(c=speed of light) (please explain)
- 0.5c,
- 0.6c,
- 0.8c,
- 0.9c,
- c
Answer #1

here,

1)

when the car stops

the final speed , v = 0 m/s

using third equation of motion

v^2 - u^2 =2 * a * D

D = - u^2 /(2a)

so, the stopping potential is directly proportional to the initial velocity

when the initial speed is doubled

the new stopping distance = (2)^2 * D = 4 D

the new stopping distance is (4D)